OpenStudy (anonymous):

if 3x^2-2x+7=0, then (x-1/3)^2=?

OpenStudy (anonymous):

you have two solutions, both complex and both with a real part of \(\frac{1}{3}\) so when you subtract you are only left with the imaginary part

OpenStudy (anonymous):

solutions are \(x=\frac{1}{3}\pm\frac{2\sqrt{5}}{3}i\) subtract and get \[\frac{2\sqrt{5}}{3}i\] then square